OpenStudy (jonathanvaldes98):

y = -3/2x + 4 Describe how you would graph this line using the slope-intercept method. i dont want the answer but how to do it.

OpenStudy (anonymous):

First give a point that x & y have it : Y = -3/2x+4 For do that do this : get x = 0 then y = 4 Now : 4 = -3/2x+4 so x = 0 draw it and then draw a line from 4 .

OpenStudy (anonymous):



OpenStudy (jonathanvaldes98):

umm... i feel like if im trying to read french... ill ask my teacher but still thanks for trying im just terrable at math tho